Mrs Pritchard has a son and a daughter. The woman is twice as old as her son and the son is twice as old as the daughter. Their ages total 91 years. how old is each person?
Answer by ankor@dixie-net.com(22740)

Mrs Pritchard has a son and a daughter.
The woman is twice as old as her son
p = 2s
and the son is twice as old as the daughter.
s = 2d
or divide equation by 2 and you have
d = .5s
Their ages total 91 years. how old is each person?
p + s + d = 91
Replace p and d
2s + s + .5 s = 91
3.5s = 91
s = 91/3.5
s = 26 yr is the son's age
:
I'll let you find p and d, see that the 3 ages add up to 91